# How Temporal Unrolling Supports Neural Physics Simulators

### v1

@@ Appendix E Evaluation Data
The differences in the studied training modalities are statistically significant for smaller network setups. As model sizes increase, the distribution of trained models becomes more similar. Our heavily overparameterized networks (e.g. 1.0M parameters for 48 degree of freedom KS system) all show highly accurate and stable predictions (see divergence time metric, Figure 20 Appendix C). Based on the theoretical considerations above, this means that the attractor of the learned dynamics is closely aligned with the ground truth system. These setups thus display less data shift and the benefits of unrolling are reduced. When these conditions in the overparameterized regime are met, NOG models are at a disadvantage due to their mismatch between gradients and loss landscape. However, these heavily oversized architectures are not practically relevant for scientific computing due to their weak scaling compared to numerical approaches. This evaluation confirms that for relevant, small to medium sized networks our results are statistically significant and hence that conclusions can be drawn.
@@
  • Figure 4 visualizes tables 5, 7, 8
  • Figure 8 visualizes tables 14, 15, 16
@@
  • Figure 11 visualizes tables 9, 10, 11
  • Figure 13 visualizes tables 18, 19constant
@@
Table 5: Correction GCN $\mathcal{L}_{2}$ errors on the KS system
@@
0.5k* 2.67063 2.07773 1.32860 0.61222 0.62611 0.09505 0.00220 0.00000
0.5k 3.42429 2.01830 1.15905 2.00918 0.81870 0.24652 0.00310 0.00001
2k 2.05080 1.53242 0.77671 1.58849 1.07397 0.20318 0.11706 0.00051
33k 0.42879 0.39406 0.33274 0.10139 0.11152 0.10516 0.15465 0.00278
0.2M 0.18810 0.20359 0.16093 0.04164 0.04815 0.05224 0.14164 0.03843
1M 0.12671 0.16176 0.12054 0.01942 0.03265 0.02622 0.00010 0.20174
Table 6: Correction CNN $\mathcal{L}_{2}$ errors on the KS system
@@
0.5k* 2.79276 1.82674 1.36738 0.53627 0.33042 0.18318 0.00000 0.00000
0.5k 2.04899 1.77093 1.32922 0.76169 0.47778 0.46564 0.08736 0.00045
2k 1.04701 0.97612 0.80260 0.16266 0.10389 0.08582 0.05436 0.00000
33k 0.39871 0.42818 0.34545 0.08113 0.08503 0.05931 0.13464 0.01148
0.2M 0.23991 0.22823 0.19139 0.06584 0.05224 0.03468 0.26910 0.00296
1M 0.12554 0.16273 0.09885 0.01664 0.06723 0.01865 0.01067 0.00001
2M 0.09592 0.09519 0.07090 0.01241 0.01160 0.00889 0.42393 0.00000
Table 7: Correction CNN $\mathcal{L}_{2}$ errors on the WAKE system
@@
6.3k 0.12045 0.07038 0.05515 0.02353 0.01374 0.00990 0.00001 0.00000
67k 0.07641 0.04430 0.03663 0.03321 0.01106 0.00568 0.00476 0.00076
1M 0.03389 0.02575 0.02496 0.00724 0.00733 0.00717 0.01120 0.00632
Table 8: Correction CNN $\mathcal{L}_{2}$ errors on the KOLM system
@@
32k 0.04097 0.04140 0.03645 0.00833 0.00506 0.00398 0.45147 0.09404
0.1M 0.03064 0.02576 0.02504 0.00344 0.00511 0.00242 0.02090 0.00104
0.5Mk 0.01635 0.01394 0.01420 0.00172 0.00144 0.00201 0.00444 0.01878
1M 0.01046 0.00859 0.00820 0.00084 0.00110 0.00122 0.00093 0.00035
Table 9: Prediction GCN $\mathcal{L}_{2}$ errors on the KS system
@@
0.5k* 7.98357 11.63567 10.23652 2.75299 1.57462 2.53653 0.00000 0.00526
0.5k 18.06462 5.10212 9.51381 5.02841 3.35803 4.80894 0.00000 0.00000
2k 10.41816 2.89637 5.62151 5.18653 0.92519 2.53810 0.00000 0.00033
33k 4.16912 1.44542 1.77320 2.10251 0.10168 0.68453 0.00000 0.00001
0.2M 1.13015 0.91450 0.79421 0.05839 0.08480 0.07173 0.00000 0.00000
1M 0.88423 0.81924 0.71090 0.05668 0.06682 0.07325 0.00101 0.00000
Table 10: Prediction CNN $\mathcal{L}_{2}$ errors on the KS system
@@
0.5k* 19.87167 13.16589 17.09354 0.60110 2.18501 1.16960 0.00000 0.00000
0.5k 6.63006 4.37441 6.78108 3.99266 2.30527 2.03592 0.01744 0.44051
2k 2.51767 2.34324 2.73029 0.47555 0.34352 0.97282 0.09577 0.19269
33k 1.38413 1.49134 1.38093 0.18477 0.20430 0.15411 0.04493 0.47643
0.2M 1.02191 0.97444 0.87086 0.18962 0.12227 0.05520 0.17634 0.00075
1M 0.76798 0.85323 0.64513 0.06914 0.12842 0.04222 0.00638 0.00000
Table 11: Prediction CNN $\mathcal{L}_{2}$ errors on the WAKE system

### v2

@@ Appendix F Evaluation Data
The differences in the studied training modalities are statistically significant for smaller network setups. As model sizes increase, the distribution of trained models becomes more similar. Our heavily overparameterized networks (e.g. 1.0M parameters for 48 degree of freedom KS system) all show highly accurate and stable predictions (see divergence time metric). Based on the theoretical considerations above, this means that the attractor of the learned dynamics is closely aligned with the ground truth system. These setups thus display less data shift and the benefits of unrolling are reduced. When these conditions in the overparameterized regime are met, NOG models are at a disadvantage due to their mismatch between gradients and loss landscape. However, these heavily oversized architectures are not practically relevant for scientific computing due to their weak scaling compared to numerical approaches. This evaluation confirms that for relevant, small to medium-sized networks our results are statistically significant and hence conclusions can be drawn.
@@
  • Figure 4 visualizes Tables 6, 8, 9
  • Figure 11 visualizes Tables 15, 16, 17
@@
  • Figure 14 visualizes Tables 10, 11, 12
  • Figure 16 visualizes Tables 19, 20constant
@@
Table 6: Correction GCN $\mathcal{L}_{\mathrm{rel}}$ errors on the KS system
@@
0.5k* 0.13866 0.12531 0.07415 0.03573 0.04180 0.00884 0.14228 0.00000
0.5k 0.12246 0.08551 0.04125 0.04741 0.02651 0.00716 0.00212 0.00000
2k 0.25568 0.06132 0.02138 0.80377 0.08010 0.00746 0.14434 0.10011
33k 0.00913 0.00856 0.00563 0.00292 0.00418 0.00309 0.31063 0.00036
0.2M 0.00232 0.00270 0.00173 0.00065 0.00073 0.00082 0.04417 0.00788
1M 0.00130 0.00196 0.00132 0.00039 0.00048 0.00030 0.00001 0.40292
Table 7: Correction CNN $\mathcal{L}_{\mathrm{rel}}$ errors on the KS system
@@
0.5k* 0.11646 0.09048 0.05381 0.02157 0.01912 0.00893 0.00013 0.00000
0.5k 0.07308 0.06987 0.04797 0.01431 0.01148 0.01033 0.21954 0.00000
2k 0.04174 0.03547 0.02306 0.01018 0.00640 0.00432 0.01252 0.00000
33k 0.00728 0.00797 0.00582 0.00227 0.00227 0.00205 0.17330 0.01953
0.2M 0.00318 0.00255 0.00204 0.00178 0.00069 0.00045 0.07378 0.00422
1M 0.00122 0.00155 0.00080 0.00020 0.00082 0.00016 0.04649 0.00000
2M 0.00078 0.00079 0.00052 0.00015 0.00014 0.00010 0.42027 0.00000
Table 8: Correction CNN $\mathcal{L}_{\mathrm{rel}}$ errors on the WAKE system
@@
6.3k 0.00078 0.00042 0.00026 0.00019 0.00010 0.00005 0.00000 0.00000
67k 0.00020 0.00013 0.00010 0.00008 0.00005 0.00004 0.00075 0.00000
1.0M 0.00013 0.00009 0.00008 0.00005 0.00004 0.00004 0.01031 0.00122
Table 9: Correction CNN $\mathcal{L}_{\mathrm{rel}}$ errors on the KOLM system
@@
32k 0.00406 0.00407 0.00361 0.00081 0.00045 0.00037 0.48124 0.01460
0.1M 0.00302 0.00257 0.00249 0.00030 0.00049 0.00022 0.00058 0.00000
0.5M 0.00163 0.00139 0.00141 0.00016 0.00014 0.00019 0.00001 0.00015
1M 0.00102 0.00084 0.00080 0.00008 0.00010 0.00011 0.00000 0.00000
Table 10: Prediction GCN $\mathcal{L}_{\mathrm{rel}}$ errors on the KS system
@@
0.5k* 1.84653 3.05853 2.59502 0.39942 0.95652 0.64054 0.00000 0.00004
0.5k 5.96483 1.34493 2.95659 2.59318 0.45774 1.65345 0.00000 0.00005
2k 3.30063 0.81323 1.02441 1.64430 0.10700 0.33664 0.00000 0.00000
33k 0.97240 0.20468 0.22368 0.78061 0.04296 0.02868 0.00004 0.00006
0.2M 0.19699 0.09459 0.06454 0.02300 0.00875 0.01415 0.00000 0.00000
1M 0.12911 0.09182 0.05792 0.01819 0.00936 0.01514 0.00000 0.00000
Table 11: Prediction CNN $\mathcal{L}_{\mathrm{rel}}$ errors on the KS system
@@
0.5k* 3.48794 1.54234 2.15773 0.49023 0.11446 0.26813 0.00000 0.00000
0.5k 1.52333 1.27744 1.22393 0.21687 0.40071 0.17834 0.01037 0.00001
2k 0.59395 0.50286 0.50700 0.17407 0.08076 0.09128 0.02017 0.02758
33k 0.17768 0.20454 0.20458 0.03064 0.04219 0.03852 0.01340 0.00964
0.2M 0.10441 0.09865 0.07651 0.03165 0.02102 0.01385 0.25063 0.00044
1M 0.06933 0.06417 0.02924 0.01334 0.01582 0.00596 0.13601 0.00000
Table 12: Prediction CNN $\mathcal{L}_{\mathrm{rel}}$ errors on the WAKE system